# Universal relation between $C_{T}$ and the CFT Weyl anomaly

### 12 Mar 2026 v2

@@ I Introduction
Conformal field theories (CFTs) occupy a central position in modern theoretical physics. They describe quantum field theories that are invariant under the full conformal group, which extends Poincaré symmetry by including scale transformations and special conformal transformations. This result, together with enlarged symmetry, strongly constrains the structure of correlation functions, the operator spectrum, and the renormalization‑group behavior of the theory. In any dimension $d$, the local operators of a CFT organize into irreducible representations of the conformal group, labeled by their scaling dimensions and Lorentz spins. Among these, the energy–momentum tensor $T_{ab}$ plays a distinguished role: it generates spacetime symmetries and its correlation functions encode universal dynamical data. In particular, the two‑point function of $T_{ab}$ is fixed up to an overall coefficient $C_{T}$, which serves as a measure of the number of degrees of freedom of the theory and appears in a variety of physical contexts, from entanglement entropy to conformal collider bounds. In our conventions, the two-point function of the energy-momentum tensor in flat space reads
@@
where $S_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the area of the unit (d1)(d-1)-sphere, ${\cal{I}}_{ab,cd}(x)$ represents the inversion operator on symmetric traceless tensors $\mathcal{I}_{ab,cd}=(I_{ac}I_{bd}+I_{ad}I_{bc})/2-\delta_{ab}\delta_{cd}/d$, which is made out of the inversion tensor $I_{ab}=\delta_{ab}-x_{a}x_{b}/|x|^{2}$, and $|x|$ denotes the Euclidean distance between the two points (see e.g. [1]).
@@
The trace anomaly anchors RG irreversibility, fixes stress‑tensor data in CFTs, generates quantum curvature actions, controls vacuum/Casimir/Hawking effects in curved space, and shapes quantum corrections in cosmology and black‑hole physics, among many other physical applications in quantum gravity, black hole physics, inflationary cosmology, string theory and statistical mechanics (see e.g. [6]).
The connection between $C_{T}$ and the Weyl anomaly coefficients exhibits a rich and intricate structure. Although these quantities arise from different physical considerations—flat‑space correlation functions versus curved‑space anomalies—they are not independent. In holographic theories, for example, both can be computed from the same bulk gravitational action, leading to universal relations. More generally, following the lead of Osborn and Petkou [1], since the integrated trace anomaly determines the running of the CFT partition function with the mass scale $\mu$, then this ties the Weyl‑squared term in the anomaly to the normalization of the stress‑tensor two‑point function around flat space. So that in any even $d$, $C_{T}$ is fixed by the part of the Weyl anomaly that is quadratic in the metric fluctuation, everything else—higher powers of the Weyl tensor, trivial total derivatives, topological terms—does not feed into $C_{T}$. Here is how that plays out dimension by dimension:
@@
    $\langle T\rangle=-a\,E_{6}+c_{1}I_{1}+c_{2}I_{2}+c_{3}I_{3}+ttd,$
    where $I_{1}$ and $I_{2}$ are cubic in the Weyl tensor, while $I_{3}$ contains terms of the schematic form $W\nabla^{2}W+lot+ttd$. The lower-order terms in derivatives (lotlot) clearly do not contribute to the quadratic expansion around flat space and are not directly connected to $C_{T}$. Therefore, it is the central charge $c_{3}$ that relates to $C_{T}$ [1, 7]
@@
    $C_{T}=69120\,c_{10}\quad(d=8)\,.$
  • d>>2, even: In higher dimensions, the precise structure of the type-B Weyl invariants is largely unknown, but for our purposes, we can safely focus on the term quadratic in the Weyl tensor. So that, up to higher curvature terms and trivial total derivatives, we have
@@
where the lower-order terms in derivatives (lotlot) now include the Euler density and only the pointwise Weyl invariants of the type-B trace anomaly, which are cubic and higher in curvature.
Understanding these relations is essential for characterizing the space of consistent CFTs in higher dimensions and to connect flat‑space observables with geometric responses. The present work develops this connection in full generality for even‐dimensional CFTs. We follow a holographic approach and combine the long-known holographic results for $C_{T}$ [12] and Weyl anomaly [8],[9],[17] for bulk Einstein gravity, together with a recently derived Chern-Gauss-Bonnet theorem on a compact Einstein manifold that allows one to pinpoint the term quadratic in the Weyl tensor in the relation between Euler density and $Q$-curvature. We also confirm our findings through two purely CFT computations that exploit the fact that the trace anomaly determines the scaling behavior of the CFT partition function and effective action.
@@ II.2 Holographic Weyl anomaly for bulk Einstein gravity and $Q$-curvature
As an outgrowth of the IR/UV connection[19], bulk infrared (IR) divergences can be mapped to the ultraviolet (UV) divergences of the boundary theory, and thus, the Weyl anomaly of the dual CFT are determined from the bulk action. The detailed computation of this holographic Weyl anomaly was originally carried out by Henningson and Skenderis [8, 9]. This sparked interest among conformal geometers and led Graham and Zworski to discover that the integrated holographic Weyl coincides with the integrated $Q$-curvature (a central object in Conformal Geometry).
There is a particular bulk metric, a solution to the Einstein equation, that enables straightforward computations and allows us to identify the numerical coefficients in front of the $Q$-curvature. Consider a particular Poincaré-Einstein metric $\hat{g}$ with an Einstein $g_{E}$ metric at its conformal infinity[20]
@@ II.3 Weyl-squared content of the $Q$-curvature
In particular, when applied to the Pfaffian that is polynomial in the Weyl tensor, one ends up with an explicit formula for the Pfaffian in terms of pointwise conformal invariants and a pure curvature scalar power that corresponds to the $Q$-curvature.
@@
The convention for the Pfaffian is that the volume integral on a compact $d$-manifold produces $(2\pi)^{d/2}$ times the Euler characteristics. Therefore, to translate to our convention for the Euler density is necessary that
@@
Alternatively, we can also connect the $Q$-curvature with a term quadratic in the Weyl tensor by careful examination of its leading and subleading terms in curvature (see discussion around Eq.(40))
@@
When written in terms of the $C_{TT}$ coefficient of the TT-correlator in momentum space ([55])
@@
$\boxed{C_{TT}=-\frac{(-1)^{d/2}\,d}{2\,(4\pi)^{d/2}}\cdot c}~.$ (21)
@@ III Universal relation between $C_{T}$ and $c$ : CFT derivations
it is direct to iterate with (i) and lower the exponent until one reaches $d$, and then (ii) produces the Dirac delta. The outcome of this iteration results in
@@ III.0.1 Standard derivation in 4D
In 4D, we then just need to expand Weyl’s conformal gravity $W^{2}$ on the RHS to quadratic order in the metric perturbation around flat space [1]. It is enough to consider the linear variation of the Weyl tensor that is given by
@@ III.0.2 Alternative derivation in 4D
In 4D, we can take advantage of the fact that, modulo the topological Euler density, one can trade Weyl-squared conformal gravity by Laczos conformal gravity, which is pure Ricci and essentially the critical 4D $Q$-curvature, modulo a trivial total derivative, by virtue of the Gauss-Bonnet formula. This path was undertaken by Sen and Sinha [27], who expanded the Ricci tensor and scalar on Cartesian coordinates
$R_{ab}=-\frac{1}{2}\left(\partial^{c}\rho\partial_{a}h_{bc}+\partial^{c}\partial_{b}h_{ac}-\nabla^{2}h_{ab}-\partial_{a}\partial_{b}h\right)+O(h^{2})\text{ and }R=-\partial^{a}\partial^{b}h_{ab}+\nabla^{2}h+O(h^{2}),$ (31)
@@ III.0.3 Standard derivation in 6D
Here we need to expand to quadratic order in the metric fluctuation $h_{ab}$ the six-dimensional Weyl conformal gravity $W\nabla^{2}W$ around flatspace, resulting in [28]
@@ III.0.4 Alternative derivation in 6D
III.0.5 General even dimensions: the standard construction
@@
Functional derivatives produce a factor of 2 and two more factors of 2 from the definition of the energy-momentum tensor, and the quartic operator $\Delta^{T}_{ab,cd}$ on transverse traceless tensors simply produces $(\partial^{2})^{2}$ times identity. So, the missing numerical factor is simply #=d\#=d. Comparing both sides of the equality, we arrive at the CFT to the universal relation
@@
III.0.6 General even $D$: $Q$-curvature and GJMS-like operators
@@
But there is a remarkable observation following this line of reasoning: the second metric variation of the $Q$-curvature, which can be thought of as a particular conformal gravity Lagrangian admitting Einstein solutions, is just the kinetic term of the corresponding Weyl graviton when restricted to the transverse-traceless component of the metric fluctuation. This is nothing but the GJMS-like operator on symmetric transverse-traceless two-tensors that happens to factorize into shifted Lichnerowicz Laplacians on Einstein manifolds, as shown some time ago by Matsumoto [33]. In particular, in flat space, it reduces to a power of the Laplacian. It is just a matter of keeping track of the overall coefficient to be able to obtain the numerical factor that fixes the RHS of eqn.32999We are grateful to A.A.Tseytlin for discussion of this point.
@@ IV Examples
For conformal powers of the Laplacian, GJMS operators, there was a conjectured $C_{T}$ value for up to the critical one (i.e. $(\partial^{2})^{d/2}$ in flat space) due to Osborn and Stergiou [37]. In their original paper, the authors verified the results in two, four, and six dimensions. The eight-dimensional case was confirmed shortly thereafter in [38], providing some insight into general proof, based on Conformal Partial Wave expansions of four-point functions for free fields, that they sum over conformal primaries.
@@
Taking advantage of the factorization properties of the GJMS operators on Einstein manifolds, we have been able to compute the accumulated relevant heat coefficient by using the explicit result of Avramidi[39] for the $b_{8}$ diagonal heat coefficient in 8D. The boundary anomaly was also reproduced holographically from the heat coefficients in 9D of the dual bulk massive scalar [40], extending the program initiated in [41] for 5D/4D and 7D/6D. The $c$ central charge for the 8D $P_{2k}$ operators is given by the following polynomial in $k$.
@@ 44 $\partial^{8}$ conformal scalar 256-256 1344-1344
We can also add, of course, the holographic data for bulk Einstein gravity that we used in the first place to obtain the numerical relation:
@@
Actually, we can do better and consider adding to the EH Lagrangian any (local) higher curvature term, i.e. contractions of the Riemann tensor and including as well its covariant derivatives
@@
here $L$ is a length scale introduced so as to make the coupling $\alpha_{n}$ dimensionless. The (Euclidean) AdS vacuum solution of the original EH action has radius $L$, but in general we will have an effective radius for the putative AdS vacuum solution. At linear order in the couplings $\alpha_{n}$ it will exist, and in general one needs to find a positive (generically nondegenerate) root of an algebraic equation for the ratio between the parameter $L$ and the effective radius $\tilde{L}=L/\sqrt{f_{\infty}}$.
Let us follow a shortcut route [42, 43, 44] to get this algebraic equation . This is equivalent to extremizing the gravitational action with respect to the putative AdS radius as in [36]. We exploit the scale invariance of the action to write
@@
We can now apply our simple holographic recipe to determine the holographic central charge $a$ by evaluating it on a particular Poincaré-Einstein bulk metric with an Einstein metric at its conformal infinity [44, 40]. From the pure-Ricci part, we get the action at $AdS$ precisely, and from the deviations, involving the Weyl tensor, it is enough to keep the Weyl-squared term
@@
From the first term we obtain the volume anomaly, proportional to the critical $Q$-curvature $Q_{d}$, and for the quadratic $\hat{W}^{2}$ term we complete with derivatives to obtain a particular conformal invariant on Einstein manifolds $\hat{W}_{2,1}^{(d+1)}$, quadratic in the bulk Weyl tensor, as obtained by Case et al.[13], that induces the corresponding Weyl invariant $W^{(d)}_{2,1}$ on the Einstein boundary. Keeping track of the numerical coefficients, we obtain for the holographic Weyl anomaly
@@
Remarkably, using our universal relation, we have confirmed the corresponding value for $C_{T}$
@@
where we have used the on-shell condition to trade ${\cal L}(f_{\infty})$ by $\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})$ and the fact that the latter only depends on $\tilde{L}$, as opposed to ${\cal L}(f_{\infty})$ which also depends on $L$. That makes it possible to act with derivatives with respect to $f_{\infty}$. Using the present universal relation, we end up with a prescription to compute $C_{T}$ directly from the type-A central charge $a$ in Einstein-like gravities. A comparison with [47] (eqn.18 and footnote [90] therein) is also in order.

### 19 Mar 2026 v3

@@ I Introduction
Conformal field theories (CFTs) occupy a central position in modern theoretical physics. They describe quantum field theories that are invariant under the full conformal group, which extends Poincaré symmetry by including scale transformations and special conformal transformations. This enlarged symmetry strongly constrains the structure of correlation functions, the operator spectrum, and the renormalization‑group behavior of the theory. In any dimension $d$, the local operators of a CFT organize into irreducible representations of the conformal group, labeled by their scaling dimensions and Lorentz spins. Among these, the energy–momentum tensor $T_{ab}$ plays a distinguished role: it generates spacetime symmetries and its correlation functions encode universal dynamical data. In particular, the two‑point function of $T_{ab}$ is fixed up to an overall coefficient $C_{T}$, which serves as a measure of the number of degrees of freedom of the theory and appears in a variety of physical contexts, from entanglement entropy to conformal collider bounds. In our conventions, the two-point function of the energy-momentum tensor in flat space reads
@@
where $S_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the area of the unit (d1)(d-1)-sphere, ${\cal{I}}_{ab,cd}(x)$ represents the inversion operator on symmetric traceless tensors $\mathcal{I}_{ab,cd}=(I_{ac}I_{bd}+I_{ad}I_{bc})/2-\delta_{ab}\delta_{cd}/d$, which is made of the inversion tensor $I_{ab}=\delta_{ab}-x_{a}x_{b}/|x|^{2}$, and $|x|$ denotes the Euclidean distance between the two points (see, e.g., [1]).
@@
The trace anomaly anchors RG irreversibility, fixes stress‑tensor data in CFTs, generates (non-local) quantum effective actions, controls vacuum/Casimir/Hawking effects in curved space, and shapes quantum corrections in cosmology and black‑hole physics, among many other physical applications in quantum gravity, black hole physics, inflationary cosmology, string theory and statistical mechanics (see e.g. [6] for a review).
The connection between $C_{T}$ and the Weyl anomaly coefficients exhibits a rich and intricate structure. Although these quantities arise from different physical considerations – flat‐space correlation functions versus curved‑space anomalies – they are not independent. In holographic theories, for example, both can be computed from the same bulk gravitational action, leading to universal relations. More generally, following the lead of Osborn and Petkou [1], since the integrated trace anomaly determines the running of the CFT partition function with the mass scale $\mu$, then this ties the Weyl‑squared term in the anomaly to the normalization of the stress‑tensor two‑point function around flat space. So that in any even $d$, $C_{T}$ is fixed by the part of the Weyl anomaly that is quadratic in the metric fluctuation, everything else—higher powers of the Weyl tensor, trivial total derivatives, topological terms—does not feed into $C_{T}$. Here is how that plays out dimension by dimension:
@@
    $(4\pi)^{3}\langle T\rangle=-a\,E_{6}+c_{1}I_{1}+c_{2}I_{2}+c_{3}I_{3}+ttd,$
    where $I_{1}$ and $I_{2}$ are cubic in the Weyl tensor, while $I_{3}$ contains terms of the schematic form $W\nabla^{2}W+lot+ttd$. The lower-order terms in derivatives (lotlot) clearly do not contribute to the quadratic expansion around flat space and are not directly related to $C_{T}$. Therefore, it is the central charge $c_{3}$ that relates to $C_{T}$ [1, 7]
@@
    $C_{T}=69120\,c_{{}_{10}}\quad(d=8)\,.$
  • d>>2, even: In higher dimensions, the precise structure of the type-B Weyl invariants is largely unknown, but for our present purposes, we can safely focus on the term quadratic in the Weyl tensor. So that, up to higher curvature terms and trivial total derivatives, we have
@@
where the lower-order terms in derivatives (lotlot) now include the Euler density and only the pointwise Weyl invariants of the type-B trace anomaly which are cubic and higher in curvature.
Understanding these relations is essential for characterizing the space of consistent CFTs in higher dimensions and to connect flat‑space observables with geometric responses. The present work develops this connection in full generality for even‐dimensional CFTs. We follow a holographic approach and combine the long-known holographic results for $C_{T}$ [14] and Weyl anomaly [8, 9, 10] for bulk Einstein gravity, together with the recently derived Chern-Gauss-Bonnet theorem on a compact Einstein manifold that allows one to pinpoint the term quadratic in the Weyl tensor in the relation between Euler density and $Q$-curvature. We also confirm our findings through two purely CFT computations that exploit the fact that the trace anomaly determines the scaling behavior of the CFT partition function and effective action.
@@ II.2 Holographic Weyl anomaly for bulk Einstein gravity and $Q$-curvature
As an outgrowth of the IR/UV connection[19], bulk infrared (IR) divergences can be mapped to the ultraviolet (UV) divergences of the boundary theory, and thus, the Weyl anomaly of the dual CFT are determined from the bulk action. The detailed computation of this holographic Weyl anomaly was originally carried out by Henningson and Skenderis [8, 9]. This sparked interest among conformal geometers and led Graham and Zworski [10] to discover that the integrated holographic Weyl coincides with the integrated $Q$-curvature – a central object in Conformal Geometry.
There is a particular bulk metric, a solution to the Einstein equation, that enables straightforward computations and allows us to identify the numerical coefficients in front of the $Q$-curvature. Consider the particular Poincaré-Einstein metric $\hat{g}$ with an Einstein $g_{E}$ metric at its conformal infinity[20]
@@ II.3 Weyl-squared content of the $Q$-curvature
In particular, when applied to the Pfaffian that is polynomial in the Weyl tensor, one ends up with an explicit formula for the Pfaffian in terms of pointwise conformal invariants and a pure curvature scalar power that corresponds to the critical $Q$-curvature.
@@
The convention for the Pfaffian is that the volume integral on a compact $d$-manifold produces $(2\pi)^{d/2}$ times the Euler characteristics. Therefore, to translate to our convention for the Euler density we first note that
@@
Alternatively, we can also connect the $Q$-curvature with a term quadratic in the Weyl tensor by careful examination of its leading and subleading terms in curvature (see discussion around eqn.(40))
@@
When written in terms of the $C_{TT}$ coefficient of the TT-correlator in momentum space (cf. [23])
@@
$\boxed{(4\pi)^{d/2}\,C_{TT}=-\frac{(-1)^{d/2}\;d}{2}\cdot c}~.$ (21)
@@ III Universal relation between $C_{T}$ and $c$ : CFT derivations
it is direct to iterate with (i) and lower the exponent until one reaches $d$, and then (ii) produces the Dirac delta. The outcome of this iteration is
@@ III.0.1 Standard derivation in 4D
In 4D, we need only expand Weyl’s conformal gravity $W^{2}$ on the RHS to quadratic order in the metric perturbation around flat space [1]. It is enough to consider the linear variation of the Weyl tensor that is given by
@@ III.0.2 Alternative derivation in 4D
In 4D, we can take advantage of the fact that, modulo the topological Euler density, one can trade Weyl-squared conformal gravity by Lanczos conformal gravity, which is pure Ricci and essentially the critical 4D $Q$-curvature, modulo a trivial total derivative, by virtue of the Gauss-Bonnet formula. This path was undertaken by Sen and Sinha [28], who expanded the Ricci tensor and scalar on Cartesian coordinates
$R_{ab}=-\frac{1}{2}\left(\partial^{c}\rho\partial_{a}h_{bc}+\partial^{c}\partial_{b}h_{ac}-\partial^{2}h_{ab}-\partial_{a}\partial_{b}h\right)+O(h^{2})\text{ and }R=-\partial^{a}\partial^{b}h_{ab}+\partial^{2}h+O(h^{2}),$ (31)
@@ III.0.3 Standard derivation in 6D
Here we need to expand to quadratic order in the metric fluctuation $h_{ab}$ the six-dimensional Weyl conformal gravity $I_{3}=W\nabla^{2}W+lot+ttd$ around flatspace, resulting in [29]
@@ III.0.4 Alternative derivation in 6D
III.0.5 Standard derivation in general even dimensions
@@
Functional derivatives produce a factor of 2 and two more factors of 2 from the definition of the energy-momentum tensor, and the quartic operator $\Delta^{T}_{ab,cd}$ on transverse traceless tensors simply produces $(\partial^{2})^{2}$ times identity. So, the missing numerical factor is simply #=d\#=d. Comparing both sides of the equality, we arrive within the CFT to the universal relation
@@
III.0.6 Alternative derivation: $Q$-curvature and GJMS-like operators
@@
But there is a remarkable observation following this line of reasoning: the second metric variation of the $Q$-curvature, which can be thought of as a particular conformal gravity Lagrangian admitting Einstein solutions, is just the kinetic term of the corresponding Weyl graviton when restricted to the transverse-traceless component of the metric fluctuation. This is nothing but the GJMS-like operator on symmetric transverse-traceless two-tensors that happens to factorize into shifted Lichnerowicz Laplacians on Einstein manifolds, as shown some time ago by Matsumoto [34]. In particular, in flat space, it reduces to a power of the Laplacian. It is just a matter of keeping track of the overall coefficient to be able to obtain the numerical factor that fixes the RHS of eqn.(37)  [^9] We are grateful to A.A.Tseytlin for discussion of this point.
@@ IV Examples
For conformal powers of the Laplacian, GJMS operators, there was a conjectured $C_{T}$ value for up to the critical one (i.e. $(\partial^{2})^{d/2}$ in flat space) due to Osborn and Stergiou [38]. In their original paper, the authors verified the results in two, four, and six dimensions. The eight-dimensional case was confirmed shortly thereafter in [39], providing some insight into a general proof, based on Conformal Partial Wave expansions of four-point functions for free fields that they summed over conformal primaries.
@@
Taking advantage of the factorization properties of the GJMS operators on Einstein manifolds, we have been able to compute the accumulated relevant heat coefficient by using the explicit result of Avramidi[40] for the $b_{8}$ diagonal heat coefficient in 8D. The boundary anomaly was also reproduced holographically from the heat coefficients in 9D of the dual bulk massive scalar [41], extending the program initiated in [42] for 5D/4D and 7D/6D. The $c$ central charge for the 8D $P_{2k}$ operators is given by the following polynomial in $k$
@@ 44 $\partial^{8}$ conformal scalar 256-256 1344-1344
We can also add to the list of examples, of course, the holographic data for bulk Einstein gravity that we used in the first place to obtain the numerical relation
@@
Actually, we can do better and consider adding to the EH Lagrangian any (local) higher curvature term, i.e. contractions of the Riemann tensor and including its covariant derivatives as well
@@
here $L$ is a length scale introduced so as to make the coupling $\alpha_{n}$ dimensionless. The (Euclidean) AdS vacuum solution of the original EH action has radius $L$, but in general, we will have an effective radius for the putative AdS vacuum solution. At linear order in the couplings $\alpha_{n}$ it will exist, and in general one needs to find a positive (generically nondegenerate) root of an algebraic equation for the ratio between the parameter $L$ and the effective AdS radius $\tilde{L}=L/\sqrt{f_{\infty}}$.
Let us follow a shortcut route [43, 44, 45] to get this algebraic equation . This is equivalent to extremizing the gravitational action with respect to the AdS radius as in [7]. We exploit the scale invariance of the action to write
@@
We can now apply our simple holographic recipe [45, 41] to determine the holographic central charges $a$ and $c$ by evaluating the action on the particular Poincaré-Einstein bulk metric with an Einstein metric at its conformal infinity. From the pure-Ricci part, we get the action at $AdS$ precisely, and from the deviations, involving the Weyl tensor, it is enough to keep the Weyl-squared term
@@
From the first term we obtain the volume anomaly, proportional to the critical $Q$-curvature $Q_{d}$, and for the quadratic $\hat{W}^{2}$ term we complete with derivatives to obtain the particular conformal invariant on Einstein manifolds $\hat{W}_{2,1}^{(d+1)}$, quadratic in the bulk Weyl tensor, as obtained by Case et al.[13], that induces the corresponding Weyl invariant $W^{(d)}_{2,1}$ on the Einstein boundary. Keeping track of the numerical coefficients, we obtain for the holographic Weyl anomaly
@@
Remarkably, using our universal relation, we can confirm the corresponding value for $C_{T}$ reported in [47]
@@
where we have used the on-shell condition to trade ${\cal L}(f_{\infty})$ by $\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})$ and the fact that the latter only depends on $\tilde{L}$, as opposed to ${\cal L}(f_{\infty})$ which also depends on $L$. That makes it possible to act with derivatives with respect to $f_{\infty}$. Using the present universal relation, we end up with a prescription to compute $C_{T}$ directly from the type-A central charge $a$ in Einstein-like gravities. A comparison with [47] (Eq.(18) and footnote [90] therein) is also in order.